1.By considering the derivative of the function f : [−1, 1] → R defined by f(x) =\frac{2x}{x^2+1}, show thatf^{-1} exists and find \left(f\right)^' (\frac{4}{5} ).


1
Expert's answer
2021-05-28T09:59:28-0400

The function "f(x)=\dfrac{2x}{x^2+1}" is one­-to-one function on "[-1, 1]."

Hence the function "f(x)" has inverse.


"f(x)=\dfrac{2x}{x^2+1}"

"f'(x)=(\dfrac{2x}{x^2+1})'=\dfrac{2(x^2+1-2x^2)}{(x^2+1)^2}=\dfrac{2(1-x^2 )}{(x^2+1)^2}"

Inverse FunctionTheorem

Let "f(x)" be a function that is both invertible and differentiable.

Let "y=f^{-1}(x)"  be the inverse of "f(x)."

For all "x" satisfying "f'(f^{-1}(x))\not=0,"


"\dfrac{dy}{dx}=\dfrac{d}{dx}(f^{-1}(x))=(f^{-1})'(x)=\dfrac{1}{f'(f^{-1}(x))}"


"f'(\dfrac{4}{5})=\dfrac{2(1-(\dfrac{4}{5})^2)}{((\dfrac{4}{5})^2+1)^2}=\dfrac{450}{1681}"